### v1

@@ 4.4 Deepfake-Resistant Guidance
In this section, we investigate whether incorporating Deepfake-resistant guidance during DDIM sampling enhances watermark robustness against Deepfake manipulations. The gradient scale $s$ in Algorithm 2 is set to 1k. We only incorporate SimSwap [1] as the specific Deepfake model in Deepfake-resistant guidance and test the robustness of watermark across various Deepfake models such as {SimSwap, UniFace, CSCS, StarGAN, FSRT}.
@@
SimSwap [1] 5.58% 2.03% 5.98% 1.02% 8.81% 4.83% 8.37% 1.26%
UniFace [34] 0.01% 0.00% 2.17% 1.80% 0.01% 0.00% 2.10% 1.57%
CSCS [35] 0.13% 0.09% 0.54% 0.35% 0.65% 0.58% 3.85% 2.62%
StarGAN [2] 4.63% 2.43% 3.83% 2.09% 5.54% 4.23% 5.43% 3.23%
FSRT [3] 0.12% 0.09% 4.09% 3.32% 0.49% 0.31% 10.07% 8.08%
Average 2.10% 0.93% 3.32% 1.71% 3.10% 1.99% 5.96% 3.35%
PSNR ↑ 41.30 41.05 41.96 41.41 39.08 38.89 41.33 40.27
SSIM ↑ 0.978 0.975 0.977 0.972 0.969 0.967 0.973 0.963
@@ 4.5 Visualization Result
The sampled images are shown in Fig. 4, with the five rows from top to bottom representing the original image $x_{\text{co}}$, the watermarked image $x_{\text{wm}}$, the distorted image $x_{\text{dt}}$, the residual signal of $\left|\mathcal{N}(x_{\text{wm}}-x_{\text{co}})-0.5\right|$ and $\left|\mathcal{N}(x_{\text{dt}}-x_{\text{wm}})-0.5\right|$, where $\mathcal{N}(x)=(x-min(x))/(max(x)-min(x))$. The first 11 columns display the effects of benign distortions, while the remaining columns show the effects of Deepfake manipulations. For simplicity, the original image required for face swapping and reenactment, as well as the specific attributes needed for attribute editing, are omitted. The last column shows the image reconstruction by the VQGAN [21] autoencoder. It can be observed that the watermark is embedded into the facial image in an invisible manner, without affecting the visual quality of the image.
@@
This study focuses on enhancing the robustness of watermarks for source tracking against Deepfake manipulations, without introducing additional functions. Because the incorporation of extra functionalities such as semi-robust watermarking for Deepfake detection could increase the complexity of data distribution learning for the diffusion model and compromise the traceability of robust watermarking. Despite these limitations, the primary goal of this study is to advance research on the application of diffusion models for Deepfake proactive forensics, with the hope of offering new insights and approaches for future development.
@@ 5 CONClUSION
In this work, we propose DiffMark, a diffusion-based robust watermarking framework that adapts the diffusion model for generative image watermarking. We construct facial image and watermark as conditions to guide the diffusion sampling process to progressively denoise and generate watermarked image. We design a cross information fusion module for the fusion of image features and watermark. To enhance the robustness of the watermark against Deepfake manipulations, we integrate a frozen autoencoder during training phase to simulate Deepfake manipulations. We further introduce Deepfake-resistant guidance during sampling to enhance the robustness of watermark against Deepfake manipulations to some extent. Experimental results demonstrate that DiffMark achieves good invisibility and robustness of watermark, providing reliable source tracking and copyright protection for facial images against Deepfake manipulations.

### v2

@@ 4.5 Deepfake-Resistant Guidance
In this section, we investigate whether incorporating Deepfake-resistant guidance during DDIM sampling enhances watermark robustness against Deepfake manipulations. The gradient scale $s$ in Algorithm 2 is set to 1k. We only incorporate SimSwap Chen et al. [2020] as the specific Deepfake model in Deepfake-resistant guidance and test the robustness of watermark across various Deepfake models such as {SimSwap, UniFace, CSCS, StarGAN, FSRT, VQGAN}.
@@
SimSwap Chen et al. [2020] 5.58% 1.71% 5.96% 1.16% 9.03% 4.61% 8.35% 1.30%
UniFace Xu et al. [2022] 0.01% 0.01% 2.20% 1.94% 0.01% 0.00% 2.07% 1.74%
CSCS Huang et al. [2024] 0.13% 0.10% 0.56% 0.45% 0.64% 0.56% 3.82% 2.88%
StarGAN Choi et al. [2018] 4.66% 3.49% 3.82% 2.60% 6.03% 4.46% 5.45% 3.70%
FSRT Rochow et al. [2024] 0.12% 0.10% 4.05% 3.43% 0.43% 0.30% 9.93% 7.89%
VQGAN Esser et al. [2021] 0.02% 0.02% 0.02% 0.01% 0.04% 0.02% 0.01% 0.01%
Average 1.75% 0.91% 2.77% 1.60% 2.70% 1.66% 4.94% 2.92%
PSNR $\uparrow$ 41.29 40.96 41.96 41.03 39.10 38.82 41.33 40.14
SSIM $\uparrow$ 0.978 0.975 0.977 0.968 0.970 0.966 0.973 0.960
@@ 4.6 Visualization Result
The sampled images are shown in Fig. 4, with the five rows from top to bottom representing the original image $x_{\text{co}}$, the watermarked image $x_{\text{wm}}$, the distorted image $x_{\text{dt}}$, the residual signal of $\left|\mathcal{N}(x_{\text{wm}}-x_{\text{co}})-0.5\right|$ and $\left|\mathcal{N}(x_{\text{dt}}-x_{\text{wm}})-0.5\right|$, where $\mathcal{N}(x)=(x-min(x))/(max(x)-min(x))$. The first 11 columns display the effects of benign distortions, while the remaining columns show the effects of Deepfake manipulations. For simplicity, the original image required for face swapping and reenactment, as well as the specific attributes needed for attribute editing, are omitted. The last column shows the image reconstruction by the VQGAN Esser et al. [2021] autoencoder. It can be observed that the watermark is embedded into the facial image in an invisible manner, without affecting the visual quality of the image.
Figure 4: The visual quality of facial images under various typical distortions. The rows from top to bottom display: (a) the original cover image $x_{\text{co}}$, (b) the watermarked image $x_{\text{wm}}$, (c) the distorted image $x_{\text{dt}}$, (d) the normalized residual signal between $x_{\text{wm}}$ and $x_{\text{co}}$, and (e) the normalized residual signal between $x_{\text{dt}}$ and $x_{\text{wm}}$. Each column represents a distinct distortion type. All images have a size of $256\times 256$ pixels.
@@
Table 8: Quantitative Experiments on Average Watermark Embedding and Extraction Times, as well as Peak Memory Usage during the Inference Phase, per $128\times 128$ Image.
Method Embed Time (s) Extract Time (s) Peak Mem (GB)
MBRS Jia et al. [2021] 0.0072 0.0065 0.1782
CIN Ma et al. [2022] 0.0182 0.0171 0.2049
ARWGAN Huang et al. [2023] 0.0024 0.0006 0.1916
SepMark Wu et al. [2023] 0.0109 0.0168 0.6150
EditGuard Zhang et al. [2024] 0.0105 0.0084 0.1101
LampMark Wang et al. [2024] 0.0040 0.0039 0.1083
Ours 0.1320 0.0042 0.1902
Ours (guidance) 0.7502 0.0048 0.7078
In this section, we evaluate and analyze the computational efficiency and memory usage. As shown in Table 8, the watermark extraction time of our method is comparable to that of the baseline methods, as all involve a one-step extraction process. However, due to the multi-step nature of the diffusion mechanism, watermark embedding naturally takes longer compared to the baseline methods, which use a one-step embedding process. Additionally, incorporating Deepfake-resistant guidance during the sampling process further increases the time due to the extra computations involving the Deepfake model. Nevertheless, the embedding time of 0.7502 seconds remains within an acceptable range for practical use. In terms of GPU memory usage, our method is similar to others when no guidance is applied. However, enabling Deepfake-resistant guidance increases memory usage, as the Deepfake model must remain loaded throughout the sampling process. Given the advantages of our method, including the enhanced watermark robustness and invisibility through the diffusion mechanism, we will focus on improving computational efficiency in future research. Despite these limitations, the primary goal of this study is to advance the application of diffusion models in Deepfake proactive forensics, with the hope of providing new insights and approaches for future development.
@@ 5 CONClUSION
In this work, we propose DiffMark, a diffusion-based robust watermarking framework that constructs facial image and watermark as conditions to guide the diffusion sampling process to progressively denoise and generate watermarked image. We design a cross information fusion module for the fusion of image features and watermark. To enhance the robustness of the watermark against Deepfake manipulations, we integrate a pre-trained frozen autoencoder during training phase and introduce Deepfake-resistant guidance during sampling phase. Experimental results demonstrate that DiffMark achieves high watermark invisibility and robustness. Although DiffMark provides traceability and copyright protection for facial images against Deepfake manipulations, its malicious use may raise ethical concerns, particularly regarding privacy violations when applied without consent. Future work could explore the privacy-preserving techniques such as differential privacy. Moreover, malicious actors involved in Deepfake distribution often tend to remove the watermarks in facial images, thereby compromising the effectiveness of watermarks. It is important to develop more robust watermark to address such risk. Lastly, further functionality could be developed. This research introduces Deepfake-resistant guidance to improve watermark traceability, and it may also be possible to leverage adversarial gradient guidance during the diffusion sampling phase to enhance the performance of Deepfake detectors or disrupt the face forgery effects of Deepfake models.